# Prediction and Analysis

### v1

@@ 6 Conclusion
Acknowledgment. We would like to thank Geotab Inc., the City of Kingston, and NSERC for their support of this work. We would also like to thank Kate Cowperthwaite, Debaditya Shome, and all the annotators each of whom contributed to this project.
@@ B. Keyword Search for dataset creation
For creating the dataset, the following keywords were used to search and collect the videos in CycleCrash:
Bicycle accident dashcam; Bicycle crash; Bicycle crash dashcam; Bicyclist car accident; Bicyclist car collision dashcam; Bicyclist collision; Bicyclist collision compilation; Bicyclist crash; Bike accident; Bike crash; Bike crash compilation; Bike hit car dashcam; Bike hit dashcam; Bike hit and run; Car hit cyclist; Cyclist accident; Cyclist accident dashcam; Cyclist being hit; Cyclist being hit compilation; Cyclist collision dashcam; Cyclist crash dashcam; Cyclist fall down; Cyclist fault dashcam; Cyclist hit and run; Cyclist hit dashcam; Cyclist hit pedestrian; Cyclist hits bus; Cyclist hits pole; Cyclist hitting car; Cyclist hits road sign; Cyclist hits train; Cyclist near miss; Dashcam cycle accident; Dashcam cycle accident compilation; Hit cyclist;
For the videos exhibiting cyclists navigating safely in an urban traffic scenario, without depicting risky behaviour, we use the clips from driving tour videos collected using the following search keywords:
City Driving tour; Downtown Driving tour; Urban Driving tour; Urban City Driving tour.

### v2

@@ 6 Conclusion
Acknowledgment. We would like to thank Geotab Inc., the City of Kingston, and NSERC for supporting this work, as well as Kate Cowperthwaite, Debaditya Shome, and the annotators who contributed to this project.
@@ B. Keyword Search for dataset creation
For creating the dataset, the following keywords were used to search and collect the videos in CycleCrash: Bicycle accident dashcam; Bicycle crash; Bicycle crash dashcam; Bicyclist car accident; Bicyclist car collision dashcam; Bicyclist collision; Bicyclist collision compilation; Bicyclist crash; Bike accident; Bike crash; Bike crash compilation; Bike hit car dashcam; Bike hit dashcam; Bike hit and run; Car hit cyclist; Cyclist accident; Cyclist accident dashcam; Cyclist being hit; Cyclist being hit compilation; Cyclist collision dashcam; Cyclist crash dashcam; Cyclist fall down; Cyclist fault dashcam; Cyclist hit and run; Cyclist hit dashcam; Cyclist hit pedestrian; Cyclist hits bus; Cyclist hits pole; Cyclist hitting car; Cyclist hits road sign; Cyclist hits train; Cyclist near miss; Dashcam cycle accident; Dashcam cycle accident compilation; Hit cyclist.
For the videos exhibiting cyclists navigating safely in an urban traffic scenario, without depicting risky behaviour, we use the clips from driving tour videos collected using the following search keywords: City Driving tour; Downtown Driving tour; Urban Driving tour; Urban City Driving tour.